OpenStudy (anonymous):

Please be 100% sure you know this answer before you answer it. I will fan

OpenStudy (escamer):

whats the question, bro?

OpenStudy (anonymous):

^

OpenStudy (ddos_dragon):

Tell us!

OpenStudy (escamer):

I demand to know lol

OpenStudy (koalagurl1):

???

OpenStudy (anonymous):

@KoalaGurl1 @Ddos_Dragon

OpenStudy (ddos_dragon):

You have sparked my interest, there is no turning back for me now!

OpenStudy (anonymous):

It's at the top of the picture??

OpenStudy (koalagurl1):

What's the question?

OpenStudy (ddos_dragon):

There is no picture.

OpenStudy (koalagurl1):

What picture??

OpenStudy (anonymous):

Which equations have a negative solution? @Ddos_Dragon @agevd12 @escamer @KoalaGurl1

OpenStudy (anonymous):

There's no attachment?

OpenStudy (ddos_dragon):

No

OpenStudy (ddos_dragon):

Just type them in :P

OpenStudy (anonymous):

Ok

OpenStudy (koalagurl1):

Multiple choice I assume?

OpenStudy (ddos_dragon):

Most likely.

OpenStudy (koalagurl1):

Yeah. :P @Ddos_Dragon

OpenStudy (ddos_dragon):

I remember answering a question with that start - is it Which equations have a negative solution? Choose all answers that are correct. A.1/4w=-2 B.-3=x/-7 C.6=-y/5 D.-2/3z=-8 ????

OpenStudy (ddos_dragon):

On my old account :P

OpenStudy (anonymous):

A. 1/4w = -2 B. -3 = x/(-7) C. 6= -y/5 D. -2/3z = -8

OpenStudy (anonymous):

Yeah its multiple choice

OpenStudy (ddos_dragon):

A, B and D?

OpenStudy (koalagurl1):

Oh ok.

OpenStudy (anonymous):

Yes those are the ones &Ddos_Dragon

OpenStudy (anonymous):

Are you sure?

OpenStudy (ddos_dragon):

Yeah, it seems a lot of people re-ask quite a bit.

OpenStudy (anonymous):

@Ddos_Dragon

OpenStudy (koalagurl1):

Good Job! :) @Ddos_Dragon

OpenStudy (anonymous):

😂😂

OpenStudy (ddos_dragon):

Fairly Certain.

OpenStudy (anonymous):

What grades are you all in?

OpenStudy (ddos_dragon):

Fin~

OpenStudy (koalagurl1):

7th

OpenStudy (ddos_dragon):

I already completed my Education :)

OpenStudy (koalagurl1):

Cool

OpenStudy (koalagurl1):

I'm homeschooled too.

OpenStudy (anonymous):

Ok on to the last question!! @Ddos_Dragon @KoalaGurl1

OpenStudy (koalagurl1):

What grade are you in? @angelierivera

OpenStudy (ddos_dragon):

I'd guess he/she is in an elementary school. No offense.

OpenStudy (koalagurl1):

Maybe idk. @Ddos_Dragon

OpenStudy (koalagurl1):

also @angelierivera are you a boy or a girl?

OpenStudy (koalagurl1):

Hello?

OpenStudy (ddos_dragon):

...

OpenStudy (anonymous):

Oh sorry I raged you in it its another question😂😂

OpenStudy (anonymous):

Taged*

OpenStudy (anonymous):

@Ddos_Dragon @KoalaGurl1

OpenStudy (anonymous):

I'm in 7th and I'm a girl😂😂

OpenStudy (koalagurl1):

Cool :)